OpenStudy (anonymous):

cos45cos30-sin45sin30

OpenStudy (anonymous):

cos(a+b) = cos(a)cos(b) - sin(a) sin(b) Do you know that formula? Or you can do it directly?

OpenStudy (anonymous):

\[ \cos(45) =sin(45)= \frac {\sqrt 2}2\\ \sin(30)=\frac 1 2\\ \cos(30)= \frac {\sqrt 3}{2} \]

OpenStudy (anonymous):

Replace and you are done

OpenStudy (haseeb96):